OpenStudy (anonymous):

Solve the equation 3x + 4y = 16 for x.

OpenStudy (anonymous):

\[3x+4y=16\] to get x by itself, start by dividing by 4 on both sides \[3x+y=4\] subtract y on both sides \[3x=y-4\] divide by 3 on both sides \[x=\frac{ y-4 }{ 3 }\]